If temperature of a back body increases from 70C to 2870C, then the rate of energy radiation becomes:
  1. 2 times
  2. 16 times
  3. 4 times
  4. (287/7)4

Grade:12

3 Answers

Eshan
5 years ago
T_1=7^{\circ}C=280K
T_2=287^{\circ}C=560K

Rate of energy radiation\propto T^4

Hence the rate of energy radiation becomes(\dfrac{560}{280})^4=2^4=16times.
Gitanjali Rout
5 years ago
 
T_1=7^{\circ}C=280K
T_2=287^{\circ}C=560K

Rate of energy radiation\propto T^4

Hence the rate of energy radiation becomes(\dfrac{560}{280})^4=2^4=16times
so hope that you find the answer useful.
Khimraj
5 years ago
Rate of energy radiation is directly proportional to T4
where T is temperature
THen the rate of energy radiation become (560//280)4 = 16 times.
Hope it clears.